# Current address: ]College of William and Mary, Williamsburg, Virginia 23187, USA Current address: ]Thomas Jefferson National Accelerator Facility, Newport News, Virginia 23606, USA

### v3

@@ II.5 Deuteron Nuclear Model
In lieu of a full theoretical description of the scattering process on the deuteron, we use an approximation to express structure functions of the deuteron as a sum of structure functions of the proton and the neutron, convoluted with their momentum and spin distribution inside the deuteron. Here we follow the prescription of Ref. [115]. Our approach corrects for the effects of Fermi motion and the effective nucleon polarization due to the deuteron D-state contribution. At the same time, experimentally, we avoid the kinematic region where coherent and two-body breakup effects are expected to be most significant, i.e., low energy transfer. For this reason, we extract neutron (and deuteron) structure function results only in the region GeV from our data, corresponding to virtual photon energies in excess of 240 MeV even at our lowest point. This also minimizes the contribution from quasi-elastic scattering and its radiative tail to our data, so that we should be dominated by the sum of contributions from the two individual nucleons. Since our lowest bin, with a mean value of GeV2, corresponds to a resolution of fm, we can also assume that these contributions largely add incoherently, given the root mean square (RMS) separation of the two nucleons in deuterium of nearly 4 fm. The region 1.073 GeV 1.15 GeV, when needed, can be augmented by a parametrization of single-pion production off protons and neutrons from the MAID partial wave analysis [113], where multi-particle final states do not contribute and the MAID parametrization is very reliable.
@@
with [116, 117, 118, 119, 120] accounting for the nucleon depolarization caused by the deuteron D-state.
@@ III The EG4 Experiment
The CLAS EG4 experiment [121, 122] was carried out at JLab’s experimental Hall B in 2006. Data in the range GeV were collected using six beam energies and on longitudinally polarized NH3 and ND3 targets. Since the physics goal of EG4 was focused on the EFT domain, data were collected with high statistics at very low values, which enabled us to form moments down to GeV2 for the NH3 target. As a comparison, the previous CLAS spin structure experiment EG1b [102] measured down to GeV2. Therefore, the data below GeV2 were improved significantly from this experiment.
@@ III.1 The Polarized Electron Beam and Beam Monitoring
The EG4 experiment employed beams circulated between 1 and 3 times, with the beam current ranging from 1 to 3.5 nA. The Hall B beamline is instrumented to measure and monitor the beam properties, namely, position, transverse distribution (beam profile), current, polarization, and charge asymmetry [20]. Nine RF cavities grouped in three beam position monitor (BPM) sets, located 36.0, 24.6, and 8.2 m upstream from the CLAS center, measured both the beam position and current. The latter was also monitored by synchrotron light monitors (SLM) and measured absolutely by a Faraday cup positioned at the end of the beamline, 29.0 m downstream from the CLAS center. The Faraday cup consisted of a 15-cm diameter Pb cylinder of 75 radiation lengths thickness and measured the accumulated charge. Time-derivation of the accumulated charge provided the beam current. The beam profile was periodically verified by running thin wires (“harps”) through the beam and detecting scattered electrons using photomultipliers (PMTs) located 10 cm from the beamline. The harps also served to calibrate the BPMs.
The EG4 experiment used a longitudinally polarized beam. The beam helicity was flipped at 30 Hz, taking the quartet structure of either + - - + or - + + -, with the first helicity window selected from a pseudo-random sequence. The beam polarization was regularly measured by a Möller polarimeter set at the entrance of the Hall. It consisted of detectors, transport and polarizing magnets, and a 25 m thick Permendur foil (49% Fe, 49% Co, 2% Va) acting as a polarized electron target and oriented at with respect to the beamline. A pair of 120-G Helmholtz coils thermally polarized the foil longitudinally to 7.5%. For each beam polarization measurement, the foil was inserted in the beam for about 30 minutes. The scattered (recoil) electrons were transported to the detectors (two lead-glass blocks) by two quadrupoles. The Möller asymmetry was measured by detecting the two electrons in coincidence and compared to the well-known theoretical expectation to obtain the beam polarization. A 30-min measurement provided a 1% statistical precision. The beam polarization for EG4 was typically around 86% [23], determined with a 2.5% systematic uncertainty. Two of the Möller measurements were cross-checked with a Mott polarimeter set in the CEBAF injector that measured the polarization of 5 MeV electrons [123]. These direct measurements were used to check the stability of the beam polarization, while for the data analysis the product of beam and target polarization was directly determined in elastic or quasi-elastic scattering, see Sections V.1 and V.4.
@@
The Hall B beamline was not equipped with a beam energy measurement device. Instead, the beam energy was calculated from the setting of the CEBAF magnets together with the beam position [124]. The relative accuracy reached on the beam energy measurement was a few . In addition to its instrumentation, the Hall B beamline had two pairs of magnetic coils to raster the beam over the target. This was to spread the heat uniformly over the target to avoid depolarization, see next section. The raster used a spiral pattern of 1.2 cm outer diameter.
@@ III.2 The Polarized Targets
During EG4 the polarized target was placed 1.01 m upstream from the CLAS center to increase the acceptance at low by reducing the minimum angle for scattered electrons. The targets included frozen 15NH3 and 15ND3 dynamically polarized at 1 K with a 5-T field, as well as two carbon targets and an empty cell for reference. These were the same as the targets used for previous CLAS double-polarization measurements [125]. The beam was rastered to protect the target from overheating and to minimize depolarization. The target density was 0.917 and 1.056 g/cm3 for NH3 and ND3, respectively.
The structure of the target insert for EG4 is shown in Fig. 3. The target cells were either 1.0 cm or 0.5 cm in length, for systematic studies, and were referred to as the long and short targets, respectively. The insert was placed into the target “banjo”, an approximately 1-liter vessel of 1-K liquid helium at the center of a 5-T superconducting split coil magnet. A complete description of the polarized target can be found in Ref. [126].
@@ III.3 The CLAS Spectrometer and the Cherenkov Detector
The basic structure of CLAS is shown in Fig. 4. The design of the CLAS detector was based on a toroidal magnetic field that was generated by six superconducting coils arranged around the beamline. In the EG4 experiment, the torus field setting was to bend electrons away from the beamline (outbending configuration). The magnet coils naturally separated the detector into six “sectors”, each functioning as an independent magnetic spectrometer. Each sector included four sub-detectors: drift chambers (DC), Cherenkov counters (CC), time-of-flight scintillator counters (SC), and electromagnetic calorimeters (EC). The drift chambers were located before, within, and after the torus magnetic field and performed charged particle tracking, allowing for the determination of the particle momentum from the curvature of their trajectories. The other sub-detectors were located outside the magnetic field region. In order to cover the very low region with the high detector efficiency necessary for the absolute cross section measurements, a new CC was built by the INFN-Genova group and was installed in Sector 6. The main elements of this detector are shown in Fig. 5. Basic information on the new CC can be found in Ref. [23].
@@ IV Data Analysis
The analysis presented in this section lead to the extraction of both the proton and the deuteron spin structure functions , , and their moments, previously reported in Refs. [21, 22]. We will present the general method (Section IV.1), event selection criteria and kinematic corrections (Section IV.2), particle identification performance and efficiency analysis of the new CC (Section IV.3), and background (Section IV.4).
@@ IV.3 Particle Identification and Cherenkov Efficiency
Further PID cuts were applied to the number of photoelectrons in the CC signal, requiring for NH3 and for ND3 analysis. The CC signals must also pass conditions that match their timing and geometrical location with those of the EC and DC [127]. Unlike the EC, the CC efficiency due to these PID selection criteria could not be easily obtained from detector simulation, and had to be extracted from data and applied to the simulation. We discuss below the details of the CC efficiency study.
@@ V.1 Simulation of (Quasi-)Elastic Polarized Yield Differences
A simulation for elastic and quasi-elastic scattering events was performed using GSIM [128, 129], the standard CLAS simulation program, combined with an event generator for (quasi-)elastic scattering. Events were generated for ranges of and . The and ranges ensured the whole Sector 6 was covered, while keeping a reasonable generator efficiency.
Two functions were used to generate separately the proton elastic peak and its radiative tail with appropriate relative weights. The proton form factors were from Ref. [130], and the radiative tail included both internal and external radiation effects following the prescription of Refs. [131][132]. The depolarization effect of the Bremsstrahlung photons was calculated following Ref. [133].
For the quasi-elastic events, we used a convolution prescription [115] to calculate the inclusive structure function of the deuteron stemming only from scattering elastically off a moving proton or neutron. Radiative effects were included using the code RCSLACPOL [102], which was based on the treatments of Refs. [131] and [134]. For the proton contribution, we used the same form factors as for elastic scattering; for the neutron form factors, we used a continued fraction parameterization [135] for and the two-parameter Galster fit [136] for .
The simulated data were then passed through the GSIM Post Processor (GPP) [128] to account for smearing factors in the detectors, and RECSIS [137], the standard CLAS simulation program for event reconstruction. Detailed inputs to the simulation, GPP, and corrections applied in the reconstruction are given below.
@@ V.4 Comparison of Elastic Simulation with Data
The uncertainty in the elastic peak normalization included the uncertainty in the proton form factor input [138], which was found to be less than at 1.1 GeV and increased slightly for higher beam energies. The uncertainty due to radiative corrections was found to be for all energies. In addition, we compared [130] with a recent proton form factor fit [139] that accounted for the uncertainty in the proton radius, and found the effect to be less than 1%. Overall, we estimated the uncertainty in elastic normalization to be with the beam energy in GeV. For the deuteron normalization based on the quasi-elastic peak, we estimated a total uncertainty of 10% for each beam energy, accounting for form factor uncertainties, the observed variation of the measured yield to simulated yield ratio with , and uncertainties in the folding and radiative correction procedure. These uncertainties were taken into account in our estimate of the systematic uncertainties (see Section V.9).
@@ V.5 Simulation of Inelastic Polarized Yield Differences
We used an event generator originally developed for Ref. [140] and improved for EG4 to generate events distributed according to the polarized cross section model from RCSLACPOL [102]. The generator worked in two steps: first, it generated separate maps for radiated inclusive polarized cross section differences
@@ V.6 Extraction of and
For asymmetries, the latest fit performed by the EG1b collaboration [102] was used. The parameterization itself is unpublished. The asymmetry model includes three options:
The asymmetries and in the resonance region are based on a combination of the deep inelastic scattering (DIS) fit with extensions into the resonance region. A parametrized fit from MAID is used to get the resonance behavior. All JLab data as of the early 2010’s, as well as some MIT BATES and NIKHEF data, were used in determining the parametrization. Sum rules are obeyed as much as possible (such as the Burkhardt-Cottingham sum rule [143] and the Soffer inequality [144]).
The asymmetry in the DIS region was based on a fit to world data on inclusive spin structure functions in the DIS region (COMPASS, SLAC, HERMES, and JLab early 2010’s) [32, 33, 34, 35, 36, 37, 145, 146, 105, 106, 45, 46, 48, 47, 108, 50, 38, 89, 102, 98, 103, 147]. The measured values were fit, with simple Regge (powers of ) behavior assumed as . The full uncertainty matrix of the fit was used to evaluate the systematic uncertainties.
@@ V.7 Systematic Uncertainty due to Models Used In the Extraction
Variation 6: The value of was changed by adding an extra twist-3 term, obtained by fitting to the SLAC E155x data [149].
@@ V.8 Contamination from Polarized Nitrogen
To account for each contribution from the 15N background, one first needs to obtain the polarization of 15N as follows: the ratio of 15N polarization over that of the proton is about [125], and there are three protons for every 15N nucleus in NH3. Therefore, the 15N contamination to either the inelastic or elastic polarized yield can be obtained by multiplying the calculated 15N unpolarized yield by the effective polarization ratio of 15N.
The contribution from 15N nuclear elastic scattering was calculated using the 15N form factors [150], and was found to be problematic only in the low bins. This is because the invariant mass for the 15N nuclear elastic scattering, calculated using Eq. (3), is separated from that of the proton elastic peak () by:

### v4

@@ II.5 Deuteron Nuclear Model
In lieu of a full theoretical description of the scattering process on the deuteron, we use an approximation to express structure functions of the deuteron as a sum of structure functions of the proton and the neutron, convoluted with their momentum and spin distribution inside the deuteron. Here we follow the prescription of Ref. [Kahn:2008nq]. Our approach corrects for the effects of Fermi motion and the effective nucleon polarization due to the deuteron D-state contribution. At the same time, experimentally, we avoid the kinematic region where coherent and two-body breakup effects are expected to be most significant, i.e., low energy transfer. For this reason, we extract neutron (and deuteron) structure function results only in the region GeV from our data, corresponding to virtual photon energies in excess of 240 MeV even at our lowest point. This also minimizes the contribution from quasi-elastic scattering and its radiative tail to our data, so that we should be dominated by the sum of contributions from the two individual nucleons. Since our lowest bin, with a mean value of GeV2, corresponds to a resolution of fm, we can also assume that these contributions largely add incoherently, given the root mean square (RMS) separation of the two nucleons in deuterium of nearly 4 fm. The region 1.073 GeV 1.15 GeV, when needed, can be augmented by a parametrization of single-pion production off protons and neutrons from the MAID partial wave analysis [Drechsel:1998hk], where multi-particle final states do not contribute and the MAID parametrization is very reliable.
@@
with [Lacombe:1980dr, Machleidt:1987hj, Zuilhof:1980ae, Kotthoff:1976fg, Desplanques:1988mp] accounting for the nucleon depolarization caused by the deuteron D-state.
@@ III The EG4 Experiment
The CLAS EG4 experiment [PR03006, PR05111] was carried out at JLab’s experimental Hall B in 2006. Data in the range GeV were collected using six beam energies and on longitudinally polarized NH3 and ND3 targets. Since the physics goal of EG4 was focused on the EFT domain, data were collected with high statistics at very low values, which enabled us to form moments down to GeV2 for the NH3 target. As a comparison, the previous CLAS spin structure experiment EG1b [CLAS:2017qga] measured down to GeV2. Therefore, the data below GeV2 were improved significantly from this experiment.
@@ III.1 The Polarized Electron Beam and Beam Monitoring
The EG4 experiment employed beams circulated between 1 and 3 times, with the beam current ranging from 1 to 3.5 nA. The Hall B beamline is instrumented to measure and monitor the beam properties, namely, position, transverse distribution (beam profile), current, polarization, and charge asymmetry [CLAS:2003umf]. Nine RF cavities grouped in three beam position monitor (BPM) sets, located 36.0, 24.6, and 8.2 m upstream from the CLAS center, measured both the beam position and current. The latter was also monitored by synchrotron light monitors (SLM) and measured absolutely by a Faraday cup positioned at the end of the beamline, 29.0 m downstream from the CLAS center. The Faraday cup consisted of a 15-cm diameter Pb cylinder of 75 radiation lengths thickness and measured the accumulated charge. Time-derivation of the accumulated charge provided the beam current. The beam profile was periodically verified by running thin wires (“harps”) through the beam and detecting scattered electrons using photomultipliers (PMTs) located 10 cm from the beamline. The harps also served to calibrate the BPMs.
The EG4 experiment used a longitudinally polarized beam. The beam helicity was flipped at 30 Hz, taking the quartet structure of either + - - + or - + + -, with the first helicity window selected from a pseudo-random sequence. The beam polarization was regularly measured by a Möller polarimeter set at the entrance of the Hall. It consisted of detectors, transport and polarizing magnets, and a 25 m thick Permendur foil (49% Fe, 49% Co, 2% Va) acting as a polarized electron target and oriented at with respect to the beamline. A pair of 120-G Helmholtz coils thermally polarized the foil longitudinally to 7.5%. For each beam polarization measurement, the foil was inserted in the beam for about 30 minutes. The scattered (recoil) electrons were transported to the detectors (two lead-glass blocks) by two quadrupoles. The Möller asymmetry was measured by detecting the two electrons in coincidence and compared to the well-known theoretical expectation to obtain the beam polarization. A 30-min measurement provided a 1% statistical precision. The beam polarization for EG4 was typically around 86% [CLAS:2016qxj], determined with a 2.5% systematic uncertainty. Two of the Möller measurements were cross-checked with a Mott polarimeter set in the CEBAF injector that measured the polarization of 5 MeV electrons [Price:1998xd]. These direct measurements were used to check the stability of the beam polarization, while for the data analysis the product of beam and target polarization was directly determined in elastic or quasi-elastic scattering, see Sections V.1 and V.4.
@@
The Hall B beamline was not equipped with a beam energy measurement device. Instead, the beam energy was calculated from the setting of the CEBAF magnets together with the beam position [Krafft:2000nm]. The relative accuracy reached on the beam energy measurement was a few . In addition to its instrumentation, the Hall B beamline had two pairs of magnetic coils to raster the beam over the target. This was to spread the heat uniformly over the target to avoid depolarization, see next section. The raster used a spiral pattern of 1.2 cm outer diameter.
@@ III.2 The Polarized Targets
During EG4 the polarized target was placed 1.01 m upstream from the CLAS center to increase the acceptance at low by reducing the minimum angle for scattered electrons. The targets included frozen 15NH3 and 15ND3 dynamically polarized at 1 K with a 5-T field, as well as two carbon targets and an empty cell for reference. These were the same as the targets used for previous CLAS double-polarization measurements [Crabb:1995xi]. The beam was rastered to protect the target from overheating and to minimize depolarization. The target density was 0.917 and 1.056 g/cm3 for NH3 and ND3, respectively.
The structure of the target insert for EG4 is shown in Fig. 3. The target cells were either 1.0 cm or 0.5 cm in length, for systematic studies, and were referred to as the long and short targets, respectively. The insert was placed into the target “banjo”, an approximately 1-liter vessel of 1-K liquid helium at the center of a 5-T superconducting split coil magnet. A complete description of the polarized target can be found in Ref. [Keith:2003ca].
@@ III.3 The CLAS Spectrometer and the Cherenkov Detector
The basic structure of CLAS is shown in Fig. 4. The design of the CLAS detector was based on a toroidal magnetic field that was generated by six superconducting coils arranged around the beamline. In the EG4 experiment, the torus field setting was to bend electrons away from the beamline (outbending configuration). The magnet coils naturally separated the detector into six “sectors”, each functioning as an independent magnetic spectrometer. Each sector included four sub-detectors: drift chambers (DC), Cherenkov counters (CC), time-of-flight scintillator counters (SC), and electromagnetic calorimeters (EC). The drift chambers were located before, within, and after the torus magnetic field and performed charged particle tracking, allowing for the determination of the particle momentum from the curvature of their trajectories. The other sub-detectors were located outside the magnetic field region. In order to cover the very low region with the high detector efficiency necessary for the absolute cross section measurements, a new CC was built by the INFN-Genova group and was installed in Sector 6. The main elements of this detector are shown in Fig. 5. Basic information on the new CC can be found in Ref. [CLAS:2016qxj].
@@ IV Data Analysis
The analysis presented in this section lead to the extraction of both the proton and the deuteron spin structure functions , , and their moments, previously reported in Refs. [CLAS:2017ozc, CLAS:2021apd]. We will present the general method (Section IV.1), event selection criteria and kinematic corrections (Section IV.2), particle identification performance and efficiency analysis of the new CC (Section IV.3), and background (Section IV.4).
@@ IV.3 Particle Identification and Cherenkov Efficiency
Further PID cuts were applied to the number of photoelectrons in the CC signal, requiring for NH3 and for ND3 analysis. The CC signals must also pass conditions that match their timing and geometrical location with those of the EC and DC [Osipenko]. Unlike the EC, the CC efficiency due to these PID selection criteria could not be easily obtained from detector simulation, and had to be extracted from data and applied to the simulation. We discuss below the details of the CC efficiency study.
@@ V.1 Simulation of (Quasi-)Elastic Polarized Yield Differences
A simulation for elastic and quasi-elastic scattering events was performed using GSIM [GSIM:2002a, GSIM:2002b], the standard CLAS simulation program, combined with an event generator for (quasi-)elastic scattering. Events were generated for ranges of and . The and ranges ensured the whole Sector 6 was covered, while keeping a reasonable generator efficiency.
Two functions were used to generate separately the proton elastic peak and its radiative tail with appropriate relative weights. The proton form factors were from Ref. [Arrington:2007ux], and the radiative tail included both internal and external radiation effects following the prescription of Refs. [Mo:1968cg][Tsai:1973py]. The depolarization effect of the Bremsstrahlung photons was calculated following Ref. [Olsen:1959zz].
For the quasi-elastic events, we used a convolution prescription [Kahn:2008nq] to calculate the inclusive structure function of the deuteron stemming only from scattering elastically off a moving proton or neutron. Radiative effects were included using the code RCSLACPOL [CLAS:2017qga], which was based on the treatments of Refs. [Mo:1968cg] and [Kukhto:1983pv]. For the proton contribution, we used the same form factors as for elastic scattering; for the neutron form factors, we used a continued fraction parameterization [Kubon:2001rj] for and the two-parameter Galster fit [E93-038:2003ixb] for .
The simulated data were then passed through the GSIM Post Processor (GPP) [GSIM:2002a] to account for smearing factors in the detectors, and RECSIS [RECSIS:1997], the standard CLAS simulation program for event reconstruction. Detailed inputs to the simulation, GPP, and corrections applied in the reconstruction are given below.
@@ V.4 Comparison of Elastic Simulation with Data
The uncertainty in the elastic peak normalization included the uncertainty in the proton form factor input [Arrington:2003qk], which was found to be less than at 1.1 GeV and increased slightly for higher beam energies. The uncertainty due to radiative corrections was found to be for all energies. In addition, we compared [Arrington:2007ux] with a recent proton form factor fit [Griffioen:2015hta] that accounted for the uncertainty in the proton radius, and found the effect to be less than 1%. Overall, we estimated the uncertainty in elastic normalization to be with the beam energy in GeV. For the deuteron normalization based on the quasi-elastic peak, we estimated a total uncertainty of 10% for each beam energy, accounting for form factor uncertainties, the observed variation of the measured yield to simulated yield ratio with , and uncertainties in the folding and radiative correction procedure. These uncertainties were taken into account in our estimate of the systematic uncertainties (see Section V.9).
@@ V.5 Simulation of Inelastic Polarized Yield Differences
We used an event generator originally developed for Ref. [CLAS:2003iiq] and improved for EG4 to generate events distributed according to the polarized cross section model from RCSLACPOL [CLAS:2017qga]. The generator worked in two steps: first, it generated separate maps for radiated inclusive polarized cross section differences
@@ V.6 Extraction of and
The choice for and was based on the latest fit by Bosted/Christy/Kalantarians (dated 2014), interpolating to the real photon point at very low . The most up-to-date references are [Bosted:2007xd] and [Christy:2007ve].
For asymmetries, the latest fit performed by the EG1b collaboration [CLAS:2017qga] was used. The parameterization itself is unpublished. The asymmetry model includes three options:
The asymmetries and in the resonance region are based on a combination of the deep inelastic scattering (DIS) fit with extensions into the resonance region. A parametrized fit from MAID is used to get the resonance behavior. All JLab data as of the early 2010’s, as well as some MIT BATES and NIKHEF data, were used in determining the parametrization. Sum rules are obeyed as much as possible (such as the Burkhardt-Cottingham sum rule [Burkhardt:1970ti] and the Soffer inequality [Soffer:1994ww]).
The asymmetry in the DIS region was based on a fit to world data on inclusive spin structure functions in the DIS region (COMPASS, SLAC, HERMES, and JLab early 2010’s) [SpinMuon:1993gcv, SpinMuonSMC:1997voo, SpinMuonSMC:1994met, SpinMuon:1995svc, SpinMuonSMC:1997ixm, SpinMuonSMC:1997mkb, HERMES:1997hjr, HERMES:2006jyl, E142:1996thl, E143:1994vcg, E143:1995rkd, E143:1995clm, E143:1998hbs, E154:1997xfa, E155:1999pwm, E155:2000qdr, COMPASS:2010wkz, COMPASS:2017hef, CLAS:2017qga, CLAS:2008xos, CLAS:2015otq, JeffersonLabHallA:2004tea]. The measured values were fit, with simple Regge (powers of ) behavior assumed as . The full uncertainty matrix of the fit was used to evaluate the systematic uncertainties.
The asymmetry in the DIS region was determined from [Wandzura:1977qf]. To evaluate systematic uncertainties, we added an extra twist-3 term obtained by fitting to the SLAC E155x [E155:2002iec] data.
@@ V.7 Systematic Uncertainty due to Models Used In the Extraction
Variation 2: was changed by with the uncertainty of given by Refs. [Bosted:2007xd] and [Christy:2007ve];
@@
Variation 6: The value of was changed by adding an extra twist-3 term, obtained by fitting to the SLAC E155x data [E155:2002iec].
@@ V.8 Contamination from Polarized Nitrogen
To account for each contribution from the 15N background, one first needs to obtain the polarization of 15N as follows: the ratio of 15N polarization over that of the proton is about [Crabb:1995xi], and there are three protons for every 15N nucleus in NH3. Therefore, the 15N contamination to either the inelastic or elastic polarized yield can be obtained by multiplying the calculated 15N unpolarized yield by the effective polarization ratio of 15N.
The contribution from 15N nuclear elastic scattering was calculated using the 15N form factors [Dally:1970lvj], and was found to be problematic only in the low bins. This is because the invariant mass for the 15N nuclear elastic scattering, calculated using Eq. (3), is separated from that of the proton elastic peak () by: